Gucchi:

Choose the graph that correctly represents the equation 15x − 5y = 30.

Gucchi:

1 attachment
Gucchi:

1 attachment
Gucchi:

@axie

Gucchi:

@smokeybrown

Gucchi:

@tetsxpreme

Gucchi:

am i right with D? or is it C?

SmokeyBrown:

It wouldn't be graphs C or D. For both of those graphs, when X=6, Y=0; if you plug those values into the equation, you get 15*6 - 5*0 = 30, which isn't true. I think you'd want to go with graph A. For that graph, when X=2, Y=0, and when X=0, Y=-6; plug those values into the equation to get 15*2 - 5*0 = 30 and 15*0 - 5*-6 = 30, which are both true; this also works for any other value of X and Y on the line, they will always return the correct result when you plug them into the equation like that. Another way to solve the problem would be to use algebra to convert the equation into the form y=mx+b, which would make it easy to see what kind of graph you want based on the slope and y-intercept. Either way, I think graph A would be the answer you want

Joe348:

i probably can help

Joe348:

you would want to rearrange your problem

Gucchi:

@gucchi wrote:
A as in the first one here?

Gucchi:

@smokeybrown

Gucchi:

1 attachment
Gucchi:

ty i get this now

Joe348:

so -5y+15x=30

Joe348:

divide everything by -5 so that would be y=-3x-6

Joe348:

divide everything by -5 so that would be y=-3x-6